OpenStudy (anonymous):

ax^2+bx+15 is the product of two binomials with integer coefficients. What set has all the possible values of b?

OpenStudy (anonymous):

A{8,16} B{-16,-8,8,16} C{1,3,5,15} D{-15,-3,-1,1,-5,15} It would be B right?

OpenStudy (raden):

the possiblity of product of two numbers has the result is 15, in order pairs they are (-1,-15),(1,15),(-3,-5),(3,5). if you add up each order pairs, -1 + (-15) = -16 1 + 15 = 16 -3 + (-5) = -8 3 + 5 = 8 so, the possible values of b is like your choice ;)